in a statistics course the average final exam score is 78 with a standard deviation of 2. Find the test score that corresponds to a z-score of -1.7.

74.6

81.4

92

78

Answer by amarjeeth123(570)

Mean mu=78
Standard Deviation sigma=2
(x-mu)/sigma=-1.7
Plugging in the values,
(x-78)/2=-1.7
x-78=-3.4
x=78-3.4=74.6
Answer=74.6